OpenStudy (anonymous):

x^2-4/x^2-18x+81*x^2-6x-27/6x-12

OpenStudy (mathteacher1729):

Do you know what a perfect square binomial is? Do you know how to cancel common factors in fractions? Do you know how to obtain a least common denominator? Without knowing these, I don't really know where to begin...

OpenStudy (anonymous):

\[\frac{1}{6}\left((82 x-188)+\frac{349}{(x+2)}\right)^{-1} \]